OpenStudy (anonymous):

2 f(x) = 5 – 3x² x = -1 x = 2 please show steps so i can work threw the real problem to get my answer

OpenStudy (anonymous):

You may have some wrong.

OpenStudy (anonymous):

this is what the problem is

OpenStudy (anonymous):

the"2 f(x)"may need be checked

OpenStudy (anonymous):

i took it right from my page

OpenStudy (anonymous):

f(x) = 5 – 3x² x = -1 x = 2

OpenStudy (anonymous):

may be need f(-1)=5-2=3; f(2)=5-3*4=-7

OpenStudy (anonymous):

i am not sure i do know its two diffenrt problems but the same formula i recall her saying

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ty